A cord is stretched with tension 102 N between two supports 28.4 m apart. If it takes a pulse 0.430 s to travel from one support to the other, what is the mass of the string?
Answer #1

On a string the wave propagates with velocity

v=\lambda\nu=\frac{\lambda}{T}=66m/s

where the wavelength is the chord length. This propagation velocity depends upon the string tension and linear mass density.

v=\sqrt{\frac{T}{\mu}}

And the linear mass density is

\mu=\frac{T}{v^2}=0.023 kg/m

The corresponding mass is

m=\mu L=0.66 kg